OpenStudy (bananas1234):

Which does not show a direct variation between x and y? y = 4/x y = 0.5x y = 7x y= x/9

OpenStudy (legends):

I think its y= x/9

OpenStudy (bananas1234):

How did you get that?

OpenStudy (legends):

Its just common sense if you look at the rest of the answers 4/x 0.5x 7x And then its x/9

OpenStudy (legends):

x/9 is the only one that is different from the other ones

OpenStudy (bananas1234):

.........That doent really help for more problems i may have to slove later. Thanks any way.

OpenStudy (bananas1234):

@Nnesha or @Directrix could you please show me how to get this?

Nnesha (nnesha):

again direct variation u use this equation for it y = kx

OpenStudy (bananas1234):

So do I plug in this time?

Nnesha (nnesha):

nope which equation are there in the form of y = kx if you solve this equation for k what you should get ?

Directrix (directrix):

direct variation means y = k x where k is some number. remember that x/9 can be written as 1/9 * x which means k= 1/9 so you have direct variation Words of @phi

OpenStudy (bananas1234):

I see

Directrix (directrix):

Question from @phi all the direct variations will be number times x which choice is not "number times x" ?

OpenStudy (bananas1234):

X/9

OpenStudy (bananas1234):

I think

OpenStudy (bananas1234):

@Legends I think i see where you were going with this.....i just did not understand. Thanks.

OpenStudy (legends):

@Bananas1234 awesome. Nice work

Directrix (directrix):

Why did you choose y = x/9. It is the same as y = (1/9) * x which is a number times x as phi said. @Bananas1234

Directrix (directrix):

y= x/9 shows direct variation. It is not the answer to this question.

OpenStudy (bananas1234):

Y = 0.5x?

Directrix (directrix):

phi says that is a number times x so it is direct variation.

Nnesha (nnesha):

remember this equation for direct variation \(\huge\color{green}{{\rm y = k}\rm x}\)

OpenStudy (bananas1234):

ummm im still confused...sorry :(

Directrix (directrix):

@Bananas1234 Which option has x in the denominator?

OpenStudy (bananas1234):

the first one

Directrix (directrix):

y = 4/x = 4 * (1/x) Is that a number (4) times just an X ? In other words, is x the same as 1/x ?

Directrix (directrix):

Is 9 the same as 1/9 ?

OpenStudy (bananas1234):

no

Directrix (directrix):

Okay.

Directrix (directrix):

y = 4/x = 4 * (1/x) Is that a number (4) times just an X ?

Directrix (directrix):

No. 4 * (1/x) is not a number times plain x. So, there is your answer.

OpenStudy (bananas1234):

I think i get it, it will just take some more practice ;). Thank you again.

Directrix (directrix):

You are welcome.